# OVA-Det: Open Vocabulary Aerial Object Detection with Image-Text Collaboration

### v2

@@ 3.3.1 Text-Guided Feature Enhancement
Figure 4: The Text-Guided Decoder in OVA-Det. By incorporating class information into object queries, TG-QE enables more focused retrieval of class-relevant features.
3.3.2 Text-Guided Query Enhancement
Similarly, we propose TG-QE, which operates on the decoder part. As shown in Fig. 4, TG-QE enriches object queries with class embeddings as prior knowledge so that, when these queries interact with multi-scale image features in the decoder, they can more effectively focus on class-relevant features and thus improve detection performance.
@@ 3.3.2 Text-Guided Query Enhancement
Method Backbone Source GZSD ZSD
mAP Recall mAP Recall
@@
RRFS[14] ResNet-101 CVPR22 10.2 1.6 2.7 19.1 5.8 8.9 2.2 14.3
ConstrastZSD[36] ResNet-101 TPAMI22 16.8 2.9 5.0 27.6 13.9 18.5 4.1 27.1
DescReg[40] ResNet-101 AAAI24 17.1 5.8 8.7 28.0 12.8 17.6 8.3 43.0
Yolo-World[4] YOLOv8-L CVPR24 18.5 3.3 5.6 41.4 15.2 22.2 7.9 37.1
OVA-Det(Ours) ResNet-50 - 23.8 4.7 7.9 60.7 28.3 38.6 10.9 66.1
Yolo-World* YOLOv8-L CVPR24 21.7 3.2 5.6 46.6 16.5 24.4 8.8 41.7
OVA-Det(Ours)* ResNet-50 - 23.1 5.3 8.6 62.5 32.7 42.9 11.7 70.0
Table 1: Comparison with the state-of-the-art under GZSD and ZSD on xView dataset, where * denotes that the dataset split redefines the base and novel classes to avoid leakage of novel classes
@@
mAP Recall mAP Recall mAP Recall mAP Recall
@@
OVA-Det* 79.7 7.1 13.1 94.9 52.7 67.8 14.3 75.9 72.0 2.2 4.3 86.5 65.3 74.4 7.8 78.8
Table 2: Comparison with the state-of-the-art under GZSD and ZSD on DIOR and DOTA dataset, where * denotes that the dataset split redefines the base and novel classes to avoid leakage of novel classes
4 Experiments
In this section, we aim to demonstrate that the proposed OVA-Det has the ability to break category limitations and explicitly demonstrate the effectiveness of our proposed components. We evaluate its zero-shot detection performance on three benchmark datasets [18, 21, 35], and assess its performance in traditional object detection on the Visdrone [10] and UAVDT [9] datasets.
4.1 Datasets and Experiment Setup
Datasets. In our evaluation of zero-shot aerial object detection, we utilize three benchmark datasets: xView [18], DIOR [21], and DOTA [35]. To ensure fairness and validity in our comparative experiments, we propose two distinct protocols for splitting the datasets into base and novel classes. In the first protocol, we follow the setup from DescReg [40] to perform the class split. However, we observed that there is an overlap between the novel categories in DIOR and the base categories in DOTA, and that DOTAv1.0 lacks annotations for many small instances. To prevent leakage of novel classes and to mitigate annotation inconsistencies, in the second protocol we adopt DOTAv1.5 and adjust the category splits in DOTA accordingly. For traditional aerial object detection, OVA-Det is evaluated using the Visdrone [10] and UAVDT [9] benchmark datasets. Due to the space limitation, the results on UAVDT and the detailed base/novel category splits for each dataset are provided in the supplementary material.
Evaluation Metrics. Following [40], we utilize standard detection metrics including mean Average Precision (mAP), Recall, Harmonic Mean (HM), and Frames Per Second (FPS) for our evaluations. The mAP and Recall are computed using an Intersection over Union (IoU) threshold of 0.5. The HM is computed to reflect the overall performance on base and novel classes. Additionally, we evaluate the proposed model under two evaluation settings: Zero-shot Detection (ZSD), which evaluates only novel classes, and Generalized Zero-shot Detection (GZSD), which assesses performance across both base and novel classes.
Implementation Details OVA-Det is implemented based on RT-DETR [46] with a ResNet50 backbone [13]. By incorporating an additional text encoder and contrastive loss, we form an end-to-end detection framework. All experiments were conducted on four 4090 GPUs, each with a batch size of 2. We set the number of object queries to 500 to accommodate the higher number of objects typically found in aerial images compared to natural scenes. Unless otherwise specified, the frames per second (FPS) are reported for a single 4090 GPU without the use of any additional acceleration techniques. Due to the space limitation, further details on the implementation are provided in the supplementary material.
4.2 Performance on Zero-shot Aerial Detection
As mentioned in Section 4.1, we adopt two protocols for splitting the base and novel classes. The second protocol is designed to avoid class leakage, and its results are marked with an asterisk (*) in the subsequent tables. For a comprehensive evaluation, we compare our approach not only with state-of-the-art aerial zero-shot detectors but also with the effective general open-vocabulary detection method YOLO-World [4].
Evaluation on xView. In Tab. 1, we compare the performance of OVA-Det with state-of-the-art methods on the xView dataset. In the ZSD setting, our method achieves significant improvements in both Recall and mAP, with absolute gains of 23.1% and 28.3% in Recall and 2.6% and 2.9% in mAP over the best-reported results under the first and second protocols respectively. In the GZSD setting, OVA-Det continues to demonstrate superior performance. Under the first protocol, the HM of Recall is 16.4% higher than the best-compared method. Although the HM of mAP is slightly lower than that of DescReg due to adopting a lighter backbone, our method still records the highest mAP in the base classes, and the Recall metric has nearly doubled. Under the second protocol, it surpasses YOLO-World by 3.0% in HM of mAP and 18.5% in HM of Recall.
Method Params ZSD FPSbs=1
DescReg(Faster-RCNN) $\geq$ 61M 15.2 11
Yolo-World-L 110M 24.8 35
OVA-Det(ResNet-50) 49M 37.2 36
Table 3: Comparison of various detection methods on the DIOR dataset. Speed is tested on a 4090 GPU with no additional acceleration. Results of DescReg are taken from its original paper.
I2T Align TG Strategy GZSD ZSD
C-WL I2T Loss TG-FE TG-QE
2.5 3.4
$\checkmark$ 5.6 7.0
$\checkmark$ $\checkmark$ 8.8 9.8
$\checkmark$ $\checkmark$ $\checkmark$ 10.4 12.3
$\checkmark$ $\checkmark$ $\checkmark$ 11.5 11.3
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 13.1 14.3
Table 4: Ablation study of various components of our proposed method on the DIOR dataset. C-WL indicates loading CLIP encoder weights, I2T Loss represents the image-to-text alignment loss. The last two columns represent mAPs under GZSD and ZSD.
Evaluation on DIOR and DOTA. In Tab. 2, we present the performance of OVA-Det on the DIOR and DOTA datasets. In the ZSD setting, OVA-Det demonstrates remarkable improvements in both Recall and mAP across both splits, the recall metric surpassing previous best results by 42.0%/55.3% on DIOR and 45.6%/59.3% on DOTA for the first/second split, respectively. Likewise, under the GZSD setting, OVA-Det achieves a higher HM of Recall than all compared approaches under both protocol splits, highlighting its strong capability for detecting novel categories. Extensive experiments across three benchmark datasets confirm that OVA-Det achieves notable performance gains in both ZSD and GZSD settings, especially in terms of Recall. These results underscore that OVA-Det offers more robust open-vocabulary detection capabilities in aerial detection.
@@ C-WL I2T Loss TG-FE TG-QE
Model Size and Latency In Tab. 3, we compare several methods on the DIOR dataset in terms of Params, ZSD mAP, and FPS. In particularly, compared to other methods, OVA-Det uses fewer parameters while achieving the highest ZSD mAP. Our model also achieves an inference speed of 36 FPS on a 4090 GPU, faster than competing methods such as DescReg and YOLO-World, and meeting the real-time detection requirements for various applications. These results highlight that OVA-Det successfully inherits the efficiency of its base architecture, making it particularly well-suited for open scenarios and aerial detection.
w/ suppress w/o suppress GZSD ZSD
TG-FE TG-QE TG-FE TG-QE
8.8 9.8
$\checkmark$ 9.3 11.7
$\checkmark$ 10.4 12.3
$\checkmark$ $\checkmark$ 11.6 12.5
$\checkmark$ $\checkmark$ 13.1 14.3
Table 5: Ablation study on background interference suppression.
4.3 Ablation Study
We conduct ablation experiments on the DIOR dataset to assess the contribution of each proposed component, as shown in Tab. 4. The baseline incorporates a text encoder and a vanilla contrastive loss for image-text alignment, which enables zero-shot detection of novel classes.
Ablation on Text-Guided Strategy. As illustrated in Tab. 4, introducing TG-FE or TG-QE individually yields notable performance improvements across both experimental settings. When both modules are employed together to form a text-guided encoder-decoder architecture, the GZSD mAP (HM) and ZSD mAP achieve 13.1% and 14.3%, respectively. These results indicate that the proposed text-guided strategy enable the model to focus on class-relevant features and enhance image-text alignment.
Ablation Study on Interference Suppression. Table 5 demonstrates the effectiveness of the interference suppression branch. Experiments show that using text information as guidance to enhance features in either the encoder or decoder is effective. Compared to the baseline, this approach yields improvements of 0.5% mAP and 1.2 % mAP, respectively. Introducing the suppression branch further improves detection accuracy. For instance, for TG-FE module, introducing suppression branch further improves mAP by 1.1 percentage points.
Visual Analysis on the Impact of the proposed TG-FE. Fig.5 visualize the impact of the TG-FE module on the features. Specifically, we visualized the correlation between visual features and the text prompt before and after the introduction of the TG-FE module. The first row corresponds to cases with larger objects, while the second row corresponds to cases with smaller objects. Comparing the results in the second and third columns, it is evident that after applying the TG-FE module, the correlation between visual features and the text prompt is significantly enhanced. In particular, in the second-row example where the ship is relatively small, the initial visual features exhibited a vague correlation spread across various parts of the image. After applying TG-FE, the ship’s features became much more prominent, and the background was effectively suppressed. As shown in the third row, if the text prompt does not match the image content, the entire image is suppressed. This result, viewed from a visualization perspective, demonstrates the effectiveness of the proposed TG-FE. TG-QE is analogous to TG-FE. One operates on the encoder, and the other on the decoder; together, they work synergistically to ensure the precise detection of OVA-Det.
4.4 Performance on Traditional Aerial Detection
Although OVA-Det is designed for open vocabulary scenarios, it can also be applied to traditional aerial object detection. To this end, we conducted experiments on Visdrone to assess its performance on traditional aerial object detection, and list the results in Tab. 6. For a fair comparison with previous work, we tested using different input resolutions. At an input size of $800\times 800$, OVA-Det achieved an mAP of 59.7%, an improvement of 0.8% compared to RT-DETR. Increasing the resolution to $1280\times 1280$ boosted the mAP by 2.4% beyond the previously best-reported result, achieving state-of-the-art performance. Overall, in the traditional aerial object detection field, the proposed OVA-Det still performs well; however, its performance advantage is not as significant as in the open vocabulary aerial detection field.
Method BackBone Resolution mAP
@@ 4.4 Performance on Traditional Aerial Detection
OVA-Det ResNet-50 800 $\times$ 800 59.7
OVA-Det ResNet-50 1280 $\times$ 1280 64.8
Table 6: Comparison with the state-of-the-art on the Visdrone validation set. ‘800 – 1200’ indicates the use of multi-scale inputs.
5 Conclusion

### v3

@@ III-C1 Image-Text Collaboration Encoder
Figure 4: The Text-Guided Decoder in RT-OVAD. By incorporating class information into object queries, TG-QE enables more focused retrieval of class-relevant features.
Visual-Guided Text Refinement. Due to varying viewpoints and shooting altitudes, aerial images often exhibit significant object scale variations. As a result, objects belonging to the same category may present drastically different visual appearances across scenes, making it challenging to align them with unified class embeddings. A straightforward solution is to enhance class embeddings with corresponding visual features. However, this approach may inevitably introduce irrelevant visual information—such as background interference—into the class representations.
To address this, we introduce a Visual-Guided Text Refinement (VG-TR) module following TG-FE. Specifically, VG-TR employs a cross-attention mechanism to fuse the enhanced visual features produced by TG-FE with the textual embeddings. Formally, given the refined visual features $F^{\prime}_{i}\in\mathbb{R}^{H\times W\times C}$ generated by the TG-FE module, and textual embeddings $T_{j}\in\mathbb{R}^{N\times D}$, the VG-TR module updates textual embeddings as follows:
$\text{Attn}=\frac{T_{j}^{(q)}{F^{\prime}_{i}}^{(k)T}}{\sqrt{d}},$ (6)
$T_{j}^{\prime}=T_{j}+\phi(\text{Attn})\cdot{F^{\prime}_{i}}^{(v)},$ (7)
where $T_{j}^{(q)}$ denotes the query projection of the textual embeddings, ${F^{\prime}_{i}}^{(k)}$ and ${F^{\prime}_{i}}^{(v)}$ represent the key and value projections of the refined visual features from TG-FE, respectively, and $\phi(\cdot)$ denotes the softmax function.
As a result, the generated text embeddings incorporate class-relevant instance features and contextual scene information, enabling the model to more effectively capture intra-class variations in appearance and scale. Furthermore, since the visual features used in VG-TR have already been refined by TG-FE—which explicitly suppresses background interference—the risk of contaminating class embeddings with irrelevant visual cues is substantially reduced. Ultimately, these refined and context-aware textual embeddings are utilized to compute visual-text similarity within the contrastive head.
III-C2 Text-Guided Query Enhancement
Similarly to TG-FE, we propose TG-QE, which operates on the decoder part. As shown in Fig. 4, TG-QE enriches object queries with class embeddings as prior knowledge so that, when these queries interact with multi-scale image features in the decoder, they can more effectively focus on class-relevant features and thus improve detection performance.
@@ III-C2 Text-Guided Query Enhancement
IV Experiments
In this section, we aim to demonstrate that the proposed RT-OVAD has the ability to break category limitations and explicitly demonstrate the effectiveness of our proposed components. To comprehensively evaluate our method, we conduct experiments on three benchmark tasks: open-vocabulary aerial object detection, zero-shot aerial detection, and traditional closed-set object detection.
IV-A Datasets and Experiment Setup
Datasets. For the evaluation of open-vocabulary aerial object detection, we follow the benchmark protocol established in [21], where the model is first pretrained on LAE-1M and then evaluated on three benchmark datasets: DIOR [1], DOTA-v2.0 [37], and LAE-80C [21].
For the evaluation of zero-shot aerial object detection, we use three widely-adopted benchmark datasets: xView [38], DIOR [1], and DOTA [37]. To ensure fairness and validity in our comparative experiments, we propose two distinct protocols for splitting the datasets into base and novel classes. In the first protocol, we follow the setup proposed in DescReg [18]. However, we observe that some novel classes in DIOR overlap with the base classes in DOTA, and that DOTA-v1.0 lacks sufficient annotations for small instances. To avoid class leakage and mitigate annotation inconsistencies, in the second protocol we adopt DOTAv1.5 and adjust the category splits in DOTA accordingly. Full details of the base/novel category splits used for zero-shot evaluation are provided in the appendix.
For conventional closed-set aerial object detection, RT-OVAD is further evaluated on two standard UAV benchmarks: VisDrone [39] and UAVDT [40].
Evaluation Metrics. Following [18, 21], we adopt standard object detection metrics for evaluation, including mean Average Precision (mAP), AP5050{50}50, Recall, Harmonic Mean (HM), and Frames Per Second (FPS). The mAP measures the average precision across multiple IoU thresholds ranging from 0.5 to 0.95 (in steps of 0.05). AP5050{50}50 and Recall are computed using an Intersection over Union (IoU) threshold of 0.5. The HM metric is used to reflect the overall performance across base and novel classes, and is computed as the harmonic mean of their respective scores. Additionally, for zero-shot detection tasks, we report results under two evaluation settings: Zero-shot Detection (ZSD), which evaluates performance on novel classes only, and Generalized Zero-shot Detection (GZSD), which assesses performance on both base and novel classes simultaneously.
Implementation Details RT-OVAD is implemented based on RT-DETR [25] with a ResNet50 backbone [41]. By incorporating an additional text encoder and contrastive loss, we form an end-to-end detection framework. All experiments were conducted on four 4090 GPUs, each with a batch size of 2. We set the number of object queries to 500 to accommodate the higher number of objects typically found in aerial images compared to natural scenes. Unless otherwise specified, the frames per second (FPS) are reported for a single 4090 GPU without the use of any additional acceleration techniques. Due to the space limitation, further details on the implementation are provided in the supplementary material.
TABLE I: Comparison with state-of-the-art open-vocabulary detectors on DIOR, DOTAv2.0, and LAE-80C benchmarks. All models are pre-trained on LAE-1M. FPS is tested on a 4090 GPU without additional acceleration. The best result within each split is shown in bold, and the second-best is underlined.
Method Source Pre-training Data Resolution Params FPS DIOR DOTAv2.0 LAE-80C
AP50 mAP mAP
YOLO-World [16] CVPR 2024 LAE-1M 800$\times$800 110M 35 78.0 33.5 13.6
Grounding DINO [15] ECCV 2024 LAE-1M 1333$\times$800 172M 11 83.6 46.0 17.7
LAE-DINO [21] AAAI 2025 LAE-1M 1333$\times$800 180M 10 85.5 46.8 20.2
RT-OVAD (Ours) LAE-1M 800$\times$800 54M 34 87.7 (+2.2) 53.8 (+7.0) 23.7 (+3.5)
TABLE II: Comparison with the state-of-the-art under GZSD and ZSD on the xView dataset, where * denotes that the dataset split redefines the base and novel classes to avoid leakage of novel classes. The best result within each split is shown in bold, and the second-best is underlined.
Method Source Backbone Resolution GZSD ZSD
AP50 Recall AP50 Recall
@@ IV-A Datasets and Experiment Setup
RRFS[42] CVPR22 ResNet–101 800$\times$800 10.2 1.6 2.7 19.1 5.8 8.9 2.2 14.3
ConstrastZSD[43] TPAMI22 ResNet–101 800$\times$800 16.8 2.9 5.0 27.6 13.9 18.5 4.1 27.1
DescReg[18] AAAI24 ResNet–101 800$\times$800 17.1 5.8 8.7 28.0 12.8 17.6 8.3 43.0
YOLO-World[16] CVPR24 YOLOv8-L 800$\times$800 18.5 3.3 5.6 41.4 15.2 22.2 7.9 37.1
Grounding DINO[15] ECCV24 Swin-T 1333$\times$800 20.3 4.8 7.7 47.0 22.1 30.0 8.7 58.6
LAE-DINO[21] AAAI25 Swin-T 1333$\times$800 20.5 4.7 7.6 46.2 31.2 37.3 7.2 58.3
RT-OVAD (Ours) ResNet–50 800$\times$800 21.1 5.3 8.5 58.9 34.9 43.8 9.8 60.1
YOLO-World* CVPR24 YOLOv8-L 800$\times$800 21.7 3.2 5.6 46.6 16.5 24.4 8.8 41.7
Grounding DINO* ECCV24 Swin-T 1333$\times$800 20.9 6.2 9.5 47.8 31.1 37.7 11.1 63.5
LAE-DINO* AAAI25 Swin-T 1333$\times$800 21.4 4.6 7.6 50.1 26.0 34.0 8.9 58.8
RT-OVAD (Ours)* ResNet–50 800$\times$800 22.4 7.0 10.6 60.1 36.7 45.5 12.3 62.5
TABLE III: Comparison with the state-of-the-art under GZSD and ZSD on the DIOR and DOTA datasets. Two evaluation protocols are reported: the upper block adopts the standard base/novel split, while the lower block (marked with *) follows a re-defined split to avoid novel-class leakage. The best value is shown in bold and the second-best is underlined.
@@
AP50 Recall AP50 Recall AP50 Recall AP50 Recall
@@
Grounding DINO* 78.4 6.0 11.2 93.9 65.3 77.0 9.8 81.0 71.2 2.8 5.4 86.4 47.8 61.6 7.5 75.0
LAE-DINO* 79.0 6.6 12.0 93.8 66.5 77.8 10.4 80.7 71.9 4.8 9.0 86.4 51.7 64.6 8.0 76.7
RT-OVAD* 78.5 7.6 13.9 94.2 64.9 76.8 14.1 81.2 72.2 2.9 5.6 85.9 66.6 74.9 9.6 76.6
IV-B Performance on Open-Vocabulary Aerial Detection
As shown in Table I, we evaluate the proposed RT-OVAD on the open-vocabulary aerial detection benchmark established by [21], achieving state-of-the-art (SOTA) performance across three benchmark datasets. Specifically, we compare our method with three mainstream open-vocabulary detectors designed for natural images (GLIP, Grounding DINO, and YOLO-World), as well as the recent aerial-specific detector LAE-DINO. The experimental results demonstrate that, despite using a lower-resolution input of $800\times 800$, our RT-OVAD still significantly outperforms previous SOTA methods, surpassing them by +2.2, +7.0, and +3.5 in terms of AP50 on the DIOR dataset, and mAP on DOTAv2.0 and LAE-80C datasets, respectively.
@@ IV-B Performance on Open-Vocabulary Aerial Detection
Model Size and Latency Table I also compares model complexity and inference speed. In particularly, compared to other methods, RT-OVAD uses fewer parameters while achieving the highest AP50/mAP performance on three benchmark datasets. RT-OVAD also achieves an inference speed of 34 FPS on an NVIDIA RTX 4090 GPU, notably faster than Grounding DINO (11 FPS), LAE-DINO (10 FPS), and comparable to YOLO-World (35 FPS), and meeting the real-time detection requirements for various applications. These results highlight that RT-OVAD successfully inherits the efficiency of its base architecture, making it particularly well-suited for open scenarios and aerial detection.
IV-C Performance on Zero-shot Aerial Detection
As mentioned in Section IV-A, we adopt two protocols for splitting the base and novel classes. The second protocol is designed to avoid class leakage, and its results are marked with an asterisk (*) in the subsequent tables. For a comprehensive evaluation, we compare our approach not only with state-of-the-art aerial zero-shot detectors but also with the effective open-vocabulary detection method YOLO-World [16], Grounding DINO [15] and LAE-DINO [21].
TABLE IV: Comparison with the state-of-the-art on the VisDrone validation set. “800–1200” indicates multi-scale inputs. The best value is shown in bold and the second-best is underlined.
Method Backbone Resolution AP50
@@ IV-C Performance on Zero-shot Aerial Detection
LAE-DINO[21] Swin-T 1333$\times$800 56.4
Grounding DINO[15] Swin-T 1333$\times$800 58.1
RT-OVAD ResNet-50 800$\times$800 60.1
RT-OVAD ResNet-50 1280$\times$1280 64.6
Evaluation on xView. In Tab. II, we compare the performance of RT-OVAD with state-of-the-art methods on the xView dataset. Under the ZSD setting, RT-OVAD achieves the highest AP5050{50}50 of 9.8% and Recall of 60.1% in the first protocol, outperforming Grounding DINO (+1.1% AP5050{50}50, +1.5% Recall), LAE-DINO (+2.6% AP5050{50}50, +1.8% Recall), and YOLO-World (+1.9% AP5050{50}50, +23.0% Recall). In the second protocol, RT-OVAD achieves an AP50 of 12.3%, surpassing all other compared methods.
In the GZSD setting, RT-OVAD consistently outperforms all compared methods in Recall across both base and novel classes. It achieves a novel-class AP5050{50}50 of 7.0% and Recall of 36.7% in the second protocol, achieving the highest HM in both AP and Recall.
Evaluation on DIOR and DOTA. We further evaluate RT-OVAD on the DIOR and DOTA datasets under both ZSD and GZSD settings, as shown in Tab. III. On DIOR, RT-OVAD* achieves the highest ZSD AP50 of 14.1%, surpassing LAE-DINO* and Grounding DINO* by 3.7% and 4.3%, respectively. In terms of Recall, RT-OVAD* reaches 81.2%, outperforming all compared methods and demonstrating strong generalization to novel classes. In the GZSD setting, RT-OVAD also achieves the highest harmonic mean (HM) of AP5050{50}50 at 13.9%, representing a +1.9% improvement over LAE-DINO*. On the DOTA dataset, RT-OVAD* achieves the best ZSD AP5050{50}50 of 9.6% and competitive Recall. While RT-OVAD exhibits slightly lower performance than LAE-DINO in a few GZSD metrics, it demonstrates clear advantages in model efficiency: as shown in Tab. I, RT-OVAD significantly reduces parameter count and inference latency compared to both Grounding DINO and LAE-DINO. These results underscore that RT-OVAD offers more robust open-vocabulary detection capabilities in aerial detection.
IV-D Performance on Traditional Aerial Detection
Performance on Visdrone dataset. Although RT-OVAD is designed for open vocabulary scenarios, it can also be applied to traditional aerial object detection. To this end, we conducted experiments on Visdrone to assess its performance on traditional aerial object detection, and list the results in Tab. IV. For a fair comparison with previous work, we tested using different input resolutions. At an input size of $800\times 800$, RT-OVAD achieved an mAP of 60.1%, an improvement of 1.2% compared to RT-DETR. Increasing the resolution to $1280\times 1280$ boosted the mAP by 2.2% beyond the previously best-reported result, achieving state-of-the-art performance.
Performance on UAVDT dataset. UAVDT [40] is also a benchmark for evaluation in aerial object detection. The dataset comprises 23,258 training images and 15,069 testing images with a resolution of $1024\times 540$, covering 3 classes. Besides Visdrone, to further demonstrate the performance of RT-OVAD on traditional aerial detection tasks, we evaluate our method on the UAVDT [40] dataset. As shown in Table V, RT-OVAD achieves the highest mAP performance, surpassing the best-compared method by 1.0% in mAP. In summary, our proposed RT-OVAD not only adapts to open-vocabulary aerial object detection scenarios but also retains excellent performance in closed settings.
IV-E Ablation Study
We conduct ablation experiments on the DIOR dataset to assess the contribution of each proposed component, as shown in Tab. VI. The baseline incorporates a text encoder and a vanilla contrastive loss for image-text alignment, which enables zero-shot detection of novel classes.
TABLE V: Comparison with the state-of-the-art on the UAVDT dataset. The best value is shown in bold and the second-best is underlined.
Method Backbone Resolution AP50
DMNet[6] ResNet-50 1000$\times$600 24.6
ClusDet[5] ResNet-50 1000$\times$600 26.5
GLSAN[45] ResNet-50 1000$\times$600 28.1
ARMNet[46] ResNet-50 1500$\times$800 30.4
UFPMP-Det[7] ResNet-50 1000$\times$600 38.7
RT-DETR[25] ResNet-50 800$\times$800 39.5
GFL V1 (CEASC)[4] ResNet-18 1024$\times$540 30.9
YOLO-World[16] YOLOv8-L 800$\times$800 35.8
Grounding DINO[15] Swin-T 1333$\times$800 36.0
LAE-DINO[21] Swin-T 1333$\times$800 36.5
RT-OVAD ResNet-50 800$\times$800 40.5
Ablation on Image-Text Collaboration Strategy. As illustrated in Tab. VI, introducing TG-FE or TG-QE individually yields notable performance improvements across both experimental settings. When both modules are employed together to form a text-guided encoder-decoder architecture, the GZSD mAP (HM) and ZSD mAP achieve 13.1% and 14.3%, respectively. Furthermore, when the VG-TR is introduced on top of TG-FE and TG-QE, the model obtains the highest GZSD HM of 13.9%, validating that enriching text embeddings with context-aware visual cues further boosts image-text alignment and improves generalization to novel classes. These results highlight the complementary benefits of the TG-FE, TG-QE, and VG-TR modules, and confirm the overall effectiveness of the proposed image-text collaboration design.
TABLE VI: Ablation study of each component on the DIOR dataset. C-WL: loading CLIP encoder weights; I2T Loss: image-to-text alignment loss.
I2T Align I-T Collaboration Strategy GZSD ZSD
C-WL I2T Loss TG-FE TG-QE VG-TR HM AP50
2.5 3.4
5.6 7.0
8.8 9.8
10.4 12.3
11.5 11.3
13.1 14.3
13.9 14.1
TABLE VII: Ablation study on background‐interference suppression.
w/ suppress w/o suppress GZSD ZSD
TG-FE TG-QE TG-FE TG-QE HM AP50
8.8 9.8
9.3 11.7
10.4 12.3
11.6 12.5
13.1 14.3
Ablation Study on Interference Suppression. Table VII demonstrates the effectiveness of the interference suppression branch. Experiments show that using text information as guidance to enhance features in either the encoder or decoder is effective. Compared to the baseline, this approach yields improvements of 0.5% and 1.2% in HM of AP50, respectively. Introducing the suppression branch further improves detection accuracy. For instance, for TG-QE module, introducing suppression branch further improves the HM of AP50 by 1.5%.
Visual Analysis on the Impact of the proposed TG-FE. Fig.5 visualize the impact of the TG-FE module on the features. Specifically, we visualized the correlation between visual features and the text prompt before and after the introduction of the TG-FE module. The first row corresponds to cases with larger objects, while the second row corresponds to cases with smaller objects. Comparing the results in the second and third columns, it is evident that after applying the TG-FE module, the correlation between visual features and the text prompt is significantly enhanced. In particular, in the second-row example where the ship is relatively small, the initial visual features exhibited a vague correlation spread across various parts of the image. After applying TG-FE, the ship’s features became much more prominent, and the background was effectively suppressed. As shown in the third row, if the text prompt does not match the image content, the entire image is suppressed. This result, viewed from a visualization perspective, demonstrates the effectiveness of the proposed TG-FE. VG-TR and TG-QE is analogous to TG-FE. TG-FE and VG-TR operates on the encoder, and the TG-QE on the decoder; together, they work synergistically to ensure the precise detection of RT-OVAD.
V Conclusion